# 3DRot: 3D Rotation Augmentation for RGB-Based 3D Tasks

### v2

@@ Introduction
where $P=[u,v,1]\in\mathbf{R}^{3\times 1}$ represents a pixel location in the image, $z$ represents the depth of the 3D point in camera space, and $K$ represents camera’s intrinsics.
@@
where $P_{A}\in\mathbf{R}^{3\times 1}$ belongs to image of camera A, $P_{B}\in\mathbf{R}^{3\times 1}$ belongs to image of camera B. $H_{AB}\in\mathbf{R}^{3\times 3}$ is defined by:
@@
where $R_{AB}\in\mathbf{R}^{3\times 3}$ is defined by $T_{A}=R_{AB}\,T_{B}$, $t_{AB}\in\mathbf{R}^{3\times 1}$ is defined by $T_{A}=T_{B}+t_{AB}$. $d\in\mathbf{R}^{1}$ represents the distance from the optical center of camera $B$ to plane $S$, $n^{T}\in\mathbf{R}^{3\times 1}$ represents the normal vector of the single plane $S$ and $n^{T}\,S+d=0$. (Hartley and Zisserman 2004; Szeliski 2022)
@@
It is important to note that $P_{A}$ and $P_{B}$, while elements of $\mathbf{R}^{3\times 1}$, represent points on their respective image planes in homogeneous coordinates, typically of the form $u,v,1]^{T}$. Given that the third component of these homogeneous image coordinates is inherently 1 (after normalization), the scaling factor $\frac{z_{B}}{z_{A}}$ in Eq. (8) can be conceptualized as a normalization constant denoted by $\lambda$. This scalar $\lambda$ ensures that the resulting vector $P_{A}$ is correctly scaled to its homogeneous representation where the last component is 1. Thus, Eq. (8) can be expressed in a more general form, highlighting the projective transformation up to a scale factor:
@@
Flipping Transformation
@@
General Linear Transformation
The derivation in Eq. (12) does not hinge on $M$ being a pure reflection; any invertible linear operator $A\!\in R^{3\times 3}$ can replace $M$. The camera-to-image mapping remains a single matrix product $H_{A}=K_{A}AK_{A}^{-1}$, so 2D-3D points stay geometrically consistent under the new view.
@@
Unlike reflections or rotations, a general linear (or affine) transform distorts the unit axes unequally; therefore an axis-aligned 3D bounding box is mapped to a parallelepiped rather than a rectangular cuboid. To retain the canonical $(R,t,s)$ representation after an affine transform $A\!\in\!GL(3)$, we require that the transformed cuboid again factorizes into an orthogonal rotation, a diagonal size matrix, and a translated center. Applying $A$ to the eight corners of the original box yields with $R_{A}\!\in\!SO(3)$ and $D_{A}={diag}(d_{1},d_{2},d_{3})$ yet to be determined.
@@
We require an affine matrix $A\in R^{3\times 3}$ such that, for every orthogonal matrix $O_{1}\in\mathrm{O}(3)$, there exist an orthogonal matrix $O_{2}$ and a diagonal matrix $D$ satisfying
@@
Setting $O_{1}=I$ yields $A^{\!\top}A=D^{2}$. Now recall the uniqueness theorem of second-order isotropic tensors (Reddy 2018) : If a real symmetric tensor T satisfies $Q\,T\,Q^{\top}=T$ for all $Q\in O(3)$, then $T=\lambda I$ .
@@
Image Padding and Principal-Point Realignment
Rotating an image about the pitch or roll axes warps its footprint on the image plane; the resulting view no longer fits the original rectangular support. A naive crop or resize would either discard valid pixels or break the projective consistency implied by the intrinsic matrix.
Screen coordinates place the origin at the top-left corner $(0,0)$, whereas camera intrinsics locate the principal point $(c_{x},c_{y})$ near the image center. After rotation, points that were inside the frame can land at negative coordinates if projected with the original intrinsics. Translating the image to make all pixels positive and then cropping back to the original size therefore shifts the geometric center away from the declared principal point, corrupting downstream geometry.

### v3

@@ 3.2 Projective Transformation
where $P=[u,v,1]\in\mathbb{R}^{3\times 1}$ represents a pixel location in the image, $z$ represents the depth of the 3D point in camera space, and $K$ represents camera’s intrinsics.
@@
where $P_{A}\in\mathbb{R}^{3\times 1}$ belongs to image of camera A, $P_{B}\in\mathbb{R}^{3\times 1}$ belongs to image of camera B. $H_{AB}\in\mathbb{R}^{3\times 3}$ is defined by:
@@
where $R_{AB}\in\mathbb{R}^{3\times 3}$ is defined by $T_{A}=R_{AB}\,T_{B}$, $t_{AB}\in\mathbb{R}^{3\times 1}$ is defined by $T_{A}=T_{B}+t_{AB}$. $d\in\mathbb{R}^{1}$ represents the distance from the optical center of camera $B$ to plane $S$, $n^{T}\in\mathbb{R}^{3\times 1}$ represents the normal vector of the single plane $S$ and $n^{T}\,S+d=0$. (Hartley and Zisserman, 2004; Szeliski, 2022)
@@
It is important to note that $P_{A}$ and $P_{B}$, while elements of $\mathbb{R}^{3\times 1}$, represent points on their respective image planes in homogeneous coordinates, typically of the form $u,v,1]^{T}$. Given that the third component of these homogeneous image coordinates is inherently 1 (after normalization), the scaling factor $\frac{z_{B}}{z_{A}}$ in Eq. (4) can be conceptualized as a normalization constant denoted by $\lambda$. This scalar $\lambda$ ensures that the resulting vector $P_{A}$ is correctly scaled to its homogeneous representation where the last component is 1. Thus, Eq. (4) can be expressed in a more general form, highlighting the projective transformation up to a scale factor:
@@
3.3 Flipping Transformation
Although analogous to rotation, flipping (reflection) is geometrically distinct because it reverses chirality. In our framework we implement flips as linear operators $M$ acting on the camera frame and object pose, and show—both via a projection-matrix view and via a direct camera-space construction—that pure flips and flip–rotation compositions $R_{c}=M\,R$ preserve 2D–3D geometric consistency when applied consistently to poses and projections. Our formula directly achieves a chiral reflection in camera space: we first reflect the image around the optical center, then re-orthogonalize the camera basis (Gram–Schmidt) and enforce a right-handed frame by negating the third camera basis vector (the first encodes gravity, the second horizontal orientation, and the third is free and used to fix chirality), so the final rotation remains in $SO(3)$ and avoids chirality-flipping ambiguities during projection. The detailed derivations of how $M$ transforms 3D points, poses, and image coordinates, how to compose $M$ with a preceding rotation, and how these viewpoints lead to consistent image-plane projections are deferred to Appendix C.
3.4 General Linear Transformation
The derivation in Eq. (14) does not hinge on $M$ being a pure reflection; any invertible linear operator $A\!\in\mathbb{R}^{3\times 3}$ can replace $M$. The camera-to-image mapping remains a single matrix product $H_{A}=K_{A}AK_{A}^{-1}$, so 2D-3D points stay geometry-consistent under the new view. However, unless $A$ is a scaled rotation, an axis-aligned cuboid does not remain a cuboid under $A$, so the $(R,t,s)$ parameterization cannot be preserved for arbitrary object orientations. We give a short proof in Appendix D.
3.5 Image Padding and Principal-Point Realignment
Rotating an image about the pitch or roll axes warps its footprint on the image plane; the resulting view no longer fits the original rectangular support. A naive crop or resize would either discard valid pixels or break the projective consistency implied by the intrinsic matrix. We render the rotated view on a minimal bounding canvas whose center is set to the updated principal point, so that all valid pixels are kept and the intrinsics remain geometry-consistent. The three-step construction is detailed in Appendix E.
4 Experiments
4.1 Datasets and Metrics
We evaluate 3DRot on three representative RGB-based 3D settings: monocular 3D cuboid detection, monocular depth estimation, and LiDAR+RGB 3D detection.
SUN10 and IN10 (monocular 3D detection).
Following OmniNOCS (Krishnan et al., 2024), we adopt its SUN RGB-D slice (Song et al., 2015) and retain the ten furniture-centric categories chair, table, desk, sofa, bed, nightstand, bookcase, dresser, toilet, bathtub. To assess cross-domain generalization, we further curate an IN10 split that merges SUN RGB-D, ARKitScenes (Baruch et al., 2021) and Hypersim (Roberts et al., 2021) using the same ten categories. Following Cube R-CNN, we discard objects with invalid depth or marginal image overlap and fix boxes whose third rotation-matrix column is not aligned with gravity. We report the seven metrics used in NOCSFormer and OmniNOCS (Krishnan et al., 2024): $IoU_{3D}$, TRANS [cm], ROT [deg], SIZE [cm], 5GRAV, 10HEAD, and $mAP_{0.5}$. $IoU_{3D}$ is computed using the exact volumetric intersection between predicted and ground-truth cuboids (rather than the fast approximation in the original Cube R-CNN code); AP denotes mean Average Precision at 2D IoU 0.5, and $mAP_{0.5}$ averages AP over the ten 3D IoU thresholds $\{0.05,0.10,\dots,0.50\}$. 5GRAV is the percentage of predictions whose gravity axis is within $5^{\circ}$ of the ground truth, and 10HEAD is the percentage whose heading (yaw) error is below $10^{\circ}$.
NYU Depth v2 and SUN RGB-D (monocular depth).
For depth estimation we insert 3DRot into a BTS ResNet-50 model (Lee et al., 2021) and follow its training protocol on the official NYU Depth v2 split (Nathan Silberman and Fergus, 2012). We evaluate in-domain on the NYU Depth v2 test set and cross-dataset on the SUN RGB-D test images, using the standard single-image depth metrics from Eigen et al. (Eigen et al., 2014); we adopt their original definitions (see Appendix F).
KITTI (LiDAR+RGB 3D detection).
To study compatibility with LiDAR+RGB pipelines and standard 3D augmentations, we plug 3DRot into MVX-Net (Sindagi et al., 2019) on the KITTI 3D object detection benchmark (Geiger et al., 2013). We follow the official protocol and use the object-detection split with categories Car, Pedestrian, and Cyclist, reporting Average Precision (AP) for 3D boxes (AP3D), bird’s-eye-view boxes (APBEV), 2D boxes, and orientation similarity (AOS) at the official IoU thresholds (0.7 for Car, 0.5 for Pedestrian and Cyclist) and moderate difficulty. [^1] KITTI switched from 11 to 40 recall positions (APR40) in 2019; our experiments follow the current official setting. Our ablations focus on the moderate AP3D scores, which are the most commonly reported indicator of overall 3D performance on KITTI.
[…]
SUN10 43.21 11.60 22.91 12.93 78.03 47.32 35.70
SUN10 + 3DRot 44.51 11.19 20.93 12.45 79.89 55.53 38.11
IN10 30.39 22.97 31.34 13.62 79.71 38.01
IN10 + 3DRot 30.59 22.75 31.23 13.57 81.20 37.80
Table 2: Ablation on SUN10. All runs share the same schedule as Table 1.
[…]
BS (Baseline) 43.21 11.60 22.91 12.93 78.03 47.32 35.70
BS + 0.5 F 43.35 11.24 25.36 12.68 26.59 47.97 35.84
BS + 0.5 F(kc) 43.79 11.43 20.91 12.61 78.67 55.23 36.58
BS + 0.8 R10-5-5 44.43 11.10 21.16 12.79 78.94 54.17 38.29
BS + 0.8 R30-5-5 43.89 11.22 21.70 12.60 81.30 52.38 37.53
BS + 0.8 R10-5-5 + 0.5F(kc) 44.51 11.19 20.93 12.45 79.89 55.53 38.11
Table 3: Evaluated on the SUN10, using a category set closely aligned with Cube R-CNN.
Method Train set $IoU_{3D}\,\uparrow$
Cube R-CNN SUN RGB-D 36.2
Cube R-CNN OMNI3DIN 37.8
Ours SUN RGB-D 43.21
Ours + 3DRot SUN RGB-D 44.51
Table 4: Ablation on SUN10. We remove objects shorter than 6.25% of image height to avoid downsampling loss. Center: image padding & principal-point realignment (without it, we simply crop to the rotated footprint). Chir.: chirality preservation during flips.
[…]
Baseline 43.70 22.22
3DRot (full) $\surd$ $\surd$ 44.58 20.78 38.06
w/o Center $\times$ $\surd$ 44.43 20.81 38.06
w/o Chirality $\surd$ $\times$ 44.40 23.57 37.49
Table 5: Effect of inserting 3DRot into the BTS ResNet-50 training schedule. Both models are trained on NYU Depth v2 and evaluated on NYU Depth v2 and SUN RGB-D; the Baseline uses the default BTS augmentation recipe (including standard 2D rotation and horizontal flips), while 3DRot adds our camera-centric 3D rotation on top of this schedule.
Method NYU abs_rel $\downarrow$ NYU rmse $\downarrow$ NYU a1 $\uparrow$ SUN abs_rel $\downarrow$ SUN rmse $\downarrow$ SUN a1 $\uparrow$
Baseline 0.1783 0.5683 0.7472 0.2502 0.6738 0.6049
3DRot 0.1685 0.5670 0.7548 0.2333 0.6612 0.6120
Table 6: BTS ResNet-50 trained on NYU Depth v2 with different augmentation families, evaluated on NYU (left) and SUN RGB-D (right). proj_family applies projective augmentations to RGB and depth jointly: a horizontal flip, an in-plane 2D rotation baseline (2DRot, treating depth as a fourth image channel but without enforcing a true roll around the principal point), or our camera-centric 3DRot. color_family adds stronger photometric jitter. Projective and color augmentations can be combined in the full schedule; here we isolate them to highlight their individual effects.
NYU SUN
Family Aug. abs $\downarrow$ rmse $\downarrow$ a1 $\uparrow$ abs $\downarrow$ rmse $\downarrow$ a1 $\uparrow$
base plain 0.2105 0.6765 0.6582 0.2891 0.7587 0.5366
proj_family flip 0.2027 0.6828 0.6574 0.2704 0.7408 0.5485
proj_family 2DRot 0.2059 0.6460 0.6818 0.2835 0.7567 0.5343
proj_family 3DRot 0.1940 0.6370 0.6954 0.2684 0.7414 0.5494
color_family color 0.1879 0.6102 0.7134 0.2532 0.6867 0.5932
Table 7: MVX-Net ablations on KITTI: impact of 3DRot on class-wise moderate 3D AP. $R_{a-b-c}$ denotes 3DRot settings where $a$, $b$, and $c$ are the maximum yaw, pitch, and roll angles (in degrees), respectively.
Exp. Overall Ped. Cyc. Car
BASE 63.85 55.12 58.13 78.31
R3-0-0 62.97 56.25 56.12 76.55
R0-3-0 64.87 57.78 59.32 77.52
R0-0-3 58.58 54.98 45.56 75.21
R3-3-0 65.16 59.26 57.94 78.27
Table 8: MVX-Net ablations on KITTI: interaction between GlobalRotScaleTrans (G), RandomFlip3D (F), and 3DRot (R). Each configuration is $(g,f,r)\in\{0,1\}^{3}$. GlobalRotScaleTrans is a pure point-cloud augmentation and does not modify the 2D images.
$(G,F,R)$ 3D box BEV AOS
$(0,0,0)$ 44.78 53.91 50.08
$(1,0,0)$ 61.86 68.96 66.69
$(0,1,0)$ 51.00 59.06 54.61
$(0,0,1)$ 50.80 59.65 54.58
$(1,1,0)$ 63.62 70.53 66.63
$(0,1,1)$ 56.45 63.47 62.59
$(1,0,1)$ 61.18 68.85 64.55
$(1,1,1)$ 65.16 71.09 67.18
4.2 Implementation Details
We briefly describe how we integrate 3DRot into each backbone; full architecture, optimization, and hyperparameter details are deferred to Appendix G.
Monocular 3D detection (SUN10/IN10).
For SUN10 and IN10 we attach a Cube R-CNN head (Brazil et al., 2023) to a frozen DINO-X detector (Ren et al., 2024; Liu et al., 2024), which we use purely as an off-the-shelf 2D object recognizer. 3DRot is applied during training as a camera-centric rotation (optionally combined with a chirality-preserving horizontal flip) that is consistently propagated to the RGB image, camera intrinsics, and 3D cuboid annotations.
Monocular depth (NYU Depth v2 / SUN RGB-D).
For depth estimation we use the BTS ResNet-50 implementation from Monocular-Depth-Estimation-Toolbox (Li, 2022) and follow its default training schedule on NYU Depth v2. 3DRot is inserted as an additional geometric augmentation in the image pre-processing stage, where the same camera-centric transform is applied to the RGB image, intrinsics, and depth map. All BTS results are from our reimplementation using the official released code, rather than copied from the original BTS paper.
LiDAR+RGB 3D detection (KITTI / MVX-Net).
For LiDAR+RGB 3D detection on KITTI we adopt the MVX-Faster-RCNN configuration in MMDetection3D (Contributors, 2020), which combines a 2D image backbone with a voxelized LiDAR branch. 3DRot is implemented as an extra camera-centric operator that jointly transforms the RGB image, camera intrinsics, and projected LiDAR points and is composed with the standard GlobalRotScaleTrans and RandomFlip3D augmentations. All MVX-Net results are from our MMDetection3D reimplementation with the official configuration, not directly copied from the original paper.
4.3 Main Results
Monocular 3D detection (SUN10 / IN10)
Table 1 summarizes single-run performance on the SUN10 and IN10 splits. On SUN10, inserting 3DRot into a frozen DINO-X + Cube R-CNN pipeline improves all seven pose metrics: $IoU_{3D}$ increases from 43.21 to 44.51, ROT decreases from 22.91 to 20.93, and $mAP_{0.5}$ rises from 35.70 to 38.11, with smaller but consistent gains on the cross-domain IN10 split. Table 3 further shows that, even when trained only on SUN RGB-D, our detector with 3DRot reaches 44.51 $IoU_{3D}$, outperforming Cube R-CNN trained on either SUN RGB-D (36.2) or OMNI3DIN (37.8).
Monocular depth estimation (NYU Depth v2 / SUN RGB-D).
For depth estimation, we plug 3DRot into the BTS ResNet-50 model trained on NYU Depth v2. As shown in Table 6, within the projective-geometry family 3DRot outperforms both horizontal flipping and 2D in-plane rotation on NYU and on cross-dataset SUN RGB-D. Compared to the baseline BTS schedule (Table 5), our final 3DRot configuration further reduces NYU abs-rel from 0.1783 to 0.1685 and improves $\delta<1.25$ from 0.7472 to 0.7548, while also lowering abs-rel and rmse on SUN RGB-D.
LiDAR+RGB 3D detection (KITTI).
For LiDAR+RGB 3D detection on KITTI, we insert 3DRot into MVX-Net within MMDetection3D. Table 7 shows that a small camera-centric yaw+pitch configuration (R3-3-0) improves moderate Overall AP3D from 63.85 to 65.16 without degrading Car AP, whereas a roll-only setting (R0-0-3) noticeably hurts Cyclist and Overall AP, which is consistent with KITTI’s acquisition setup where the cameras are mounted approximately level with the ground plane, so large synthetic roll angles produce atypical viewpoints. Table 8 further indicates that 3DRot can be combined with standard scene-level augmentations such as GlobalRotScaleTrans and RandomFlip3D, with the best configuration including 3DRot remaining within a few AP points of the strongest non-3DRot setting.
Additional quantitative results and full tables are deferred to Appendix I.
4.4 Ablation Study
Ablations on monocular 3D detection.
Table 2 and Table 4 dissect the SUN10 gains. Naively adding a horizontal flip (BS + 0.5F) slightly improves $IoU_{3D}$ but severely harms pose quality (ROT rises to 25.36 and 5GRAV collapses), showing that mirroring a full $3{\times}3$ rotation matrix without chirality control is unsafe. Enforcing chirality preservation (BS + 0.5F(kc)) restores low rotation error and yields a substantial 10HEAD gain, while pure camera-centric rotations (R10-5-5, R30-5-5) improve most metrics and combining them with chirality-safe flips gives the best overall trade-off. Table 4 further shows that padding and principal-point realignment bring a small but consistent benefit, whereas disabling chirality preservation significantly degrades orientation metrics, confirming that the improvements come from geometry-consistent rotations and chirality-safe reflections rather than generic image warping.
Ablations on monocular depth estimation.
For depth estimation, Table 6 isolates projective augmentations within a fixed BTS training recipe. Within this projective family, 3DRot consistently outperforms horizontal flips and 2D in-plane rotations on both NYU Depth v2 and SUN RGB-D, indicating that explicitly updating intrinsics and camera-space rays is more effective than treating rotations as purely 2D warps. Comparing Table 5 and Table 6 further shows that these gains are additive to standard color-based augmentation: even with strong photometric jitter, inserting a small-angle 3DRot still reduces abs-rel and improves $\delta<1.25$ on both datasets.
Interaction with standard 3D augmentations in MVX-Net.
We also examine how 3DRot interacts with the standard 3D scene-level augmentations in MVX-Net, namely GlobalRotScaleTrans and RandomFlip3D (Table 8). Turning on GlobalRotScaleTrans alone $(1,0,0)$ yields a large improvement over the no-augmentation setting $(0,0,0)$, raising moderate 3D AP from 44.78 to 61.86 and AOS from 50.08 to 66.69, while adding RandomFlip3D $(1,1,0)$ gives the highest 3D AP of 63.62. Configurations that also include 3DRot $(0,0,1),(0,1,1),(1,0,1)$ obtain 3D AP between 50.80 and 61.18 with AOS in the mid-50s to mid-60s; in particular, the best 3DRot configuration $(1,0,1)$ is within about 2.5 points of $(1,1,0)$ in 3D AP and roughly 2 points in AOS. Overall, these patterns suggest that camera-centric rotations can be combined with existing 3D augmentations in MVX-Net without destabilizing detection performance, while leaving headroom for further tuning of angle ranges and scheduling.
5 Conclusion & Discussion
We have introduced 3DRot, a simple yet effective rotation-and-reflection augmentation that operates directly at the camera optical center. 3DRot does not require depth information for scene reconstruction while maintaining 2D-3D geometric consistency. Despite these benefits, 3DRot has limitations. It is most effective when camera poses are diverse; on KITTI, where cameras are roughly level with the ground plane, strong roll perturbations can hurt performance (Table 7). When intrinsics are missing they must be estimated from 2D–3D correspondences, and large intrinsic variation across a dataset may also require a normalization step, e.g., mapping all views to a shared “virtual camera” as in Cube R-CNN (Brazil et al., 2023). We discuss in Appendix H.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional NOCS Visualizations
Figure 4: NOCS (R/G/B encode X/Y/Z) maps under camera-centric rotations about the optical center. The block is a $2\times 2$ grid shown in row-major order with camera rotations: $(0^{\circ},0^{\circ},0^{\circ})$, yaw $40^{\circ}$, pitch $20^{\circ}$, and roll $20^{\circ}$. In all cases 3DRot applies the same pure-rotation homography to the RGB image and updates labels/intrinsics accordingly, so the depth and NOCS remain 2D–3D consistent while the image footprint changes.
Appendix B Camera-centric rotation of 3D points and cuboids
To propagate a virtual camera rotation to all scene geometry, we first clarify how 3D points migrate between the world frame, the original camera frame, and the rotated camera frame. This section revisits the world-to-camera transform, derives the point-wise update induced by a camera-centric rotation, and then lifts the result to the cuboid pose used in the main paper.
A point expressed in the world frame is mapped to the camera frame through
$T_{c}=R\,T_{w}+t,$ (7)
where $T_{c}\in\mathbb{R}^{3\times 1}$ denotes a 3D point in the camera coordinate system, $T_{w}\in\mathbb{R}^{3\times 1}$ denotes the same point in the world coordinate system, and $R\in\mathbb{R}^{3\times 3}$ and $t\in\mathbb{R}^{3\times 1}$ are the camera extrinsics.
After rotating the camera by a camera-centric rotation matrix $R_{c}$, the new camera coordinate system is
$T_{\text{rot-c}}=R_{c}\,T_{c}.$ (8)
Combining Eq. 7 and Eq. 8 gives
$T_{\text{rot-c}}=R_{c}\,R\,T_{w}+R_{c}\,t,$ (9)
which shows that applying a camera-centric rotation requires updating the camera extrinsics from $(R,t)$ to $(R_{c}R,\,R_{c}t)$.
The derivation above describes how any single 3D point is transferred from the original camera frame to the rotated one. We now extend this result to an entire axis-aligned cuboid whose pose is parameterized by its rotation, side lengths, and translation. Let
$T_{bb}=R_{bb}\,S_{bb}+t_{bb}$ (10)
denote the $3\times 8$ matrix of the eight cuboid corners in the original camera coordinates, where the column matrix $S_{bb}\in\mathbb{R}^{3\times 8}$ is formed by concatenating the vectorized offsets $\bigl(\pm\tfrac{W}{2},\,\pm\tfrac{H}{2},\,\pm\tfrac{L}{2}\bigr)^{\!\top}$.
Applying the camera-centric rotation in Eq. (8) to every corner gives
$T_{bb}^{\text{new}}$ $=R_{c}\bigl(R_{bb}\,S_{bb}+t_{bb}\bigr)$
$=\underbrace{(R_{c}R_{bb})}_{R_{\text{new}}}\,S_{bb}+\underbrace{R_{c}t_{bb}}_{t_{\text{new}}}.$ (11)
Hence the updated pose parameters are
$R_{\text{new}}=R_{c}R_{bb},\qquad S_{\text{new}}=S_{bb},\qquad t_{\text{new}}=R_{c}t_{bb}.$ (12)
That is, the cuboid’s orientation and translation are left-multiplied by the camera rotation $R_{c}$, whereas its side lengths remain unchanged. This update rule is used throughout our experiments when applying 3DRot to 3D bounding boxes.
Appendix C Flipping Transformation Details
@@
Appendix D Loss of Cuboid Structure Under General Affine Maps
@@
Unlike reflections or rotations, a general linear (or affine) transform distorts the unit axes unequally; therefore an axis-aligned 3D bounding box is mapped to a parallelepiped rather than a rectangular cuboid. To retain the canonical $(R,t,s)$ representation after an affine transform $A\!\in\!GL(3)$, we require that the transformed cuboid again factorizes into an orthogonal rotation, a diagonal size matrix, and a translated center. Applying $A$ to the eight corners of the original box yields $R_{A}\!\in\!SO(3)$ and $D_{A}={diag}(d_{1},d_{2},d_{3})$ yet to be determined.
@@
We require an affine matrix $A\in\mathbb{R}^{3\times 3}$ such that, for every orthogonal matrix $O_{1}\in\mathrm{O}(3)$, there exist an orthogonal matrix $O_{2}$ and a diagonal matrix $D$ satisfying
@@
Setting $O_{1}=I$ yields $A^{\!\top}A=D^{2}$. Now recall the uniqueness theorem of second-order isotropic tensors (Reddy, 2018) : If a real symmetric tensor T satisfies $Q\,T\,Q^{\top}=T$ for all $Q\in O(3)$, then $T=\lambda I$ .
@@
Appendix E Image Padding and Principal-Point Realignment